OpenStudy (anonymous):

A force of 35N is exerted on a wooden block that is 80cm long, 47cm wide, and 25 cm high. Calculate the pressure applied to the box

OpenStudy (shamim):

the area of ur given block will b\[2(80\times47+80\times25+47\times25)\]

OpenStudy (shamim):

13870\[cm ^{2}\]

OpenStudy (shamim):

\[=\frac{ 13870 }{ 100\times100 }m ^{2}\]

OpenStudy (shamim):

\[=1.387m ^{2}\]

OpenStudy (shamim):

now the pressure is\[=\frac{ force }{ area }\]

OpenStudy (shamim):

\[=\frac{ 35 }{ 13870 }N/m ^{2}\]

OpenStudy (shamim):

r u following me

OpenStudy (anonymous):

Yes, so you substitute surface area for "area"?